OpenStudy (anonymous):

2log6+log (1/3)=log12 which properties of logarithmics are/is used? power property and product property quotient property and product property quotient property only power property only teacher did not cover this, but still put it on th EOC

OpenStudy (anonymous):

power: 2log6=log6^2=log36 product log36+log1/3=log36*1/3=log12

OpenStudy (anonymous):

so its both?

OpenStudy (anonymous):

yeah it is